OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a14 = -33, a15 = 9

OpenStudy (anonymous):

an = -579 + 42(n + 1) an = -579 + 42(n - 1) an = -579 - 42(n + 1) an = -579 - 42(n - 1)

OpenStudy (anonymous):

a(n) = a1 + d(n-1) where a(n) is the nth term d = common difference n = index

OpenStudy (anonymous):

so i just plug in the values?

OpenStudy (cggurumanjunath):

b

OpenStudy (anonymous):

lol thanks

OpenStudy (cggurumanjunath):

haha . Anytime .

OpenStudy (cggurumanjunath):

:) . welcome